find the exact value of cos(a+b) if sin a =4/5 and sin b= -5/13 ( - in 5 only)with a in quadrant II and b in Quadrant III
cos(a+b) = ?

Answer by stanbon(75887)

find the exact value of cos(a+b) if sin a =4/5 and sin b= -5/13 ( - in 5 only)with a in quadrant II and b in Quadrant III
cos(a+b) = cos(a)cos(b)-sin(a)sin(b)
Since sin(a) = 4/5 in QII, cos(a) = -3/5
Since sin(b) = -5/13 in QIII, cos(b) = -12/13
--------
Substitute those values and solve for cos(a+b)
---------
Cheers,
Stan H.
----------------